### v2

# Scaled Block Vecchia Approximation for High-Dimensional Gaussian Process Emulation on GPUs

Abstract.
Emulating computationally intensive scientific simulations is crucial for enabling uncertainty quantification, optimization, and informed decision-making at scale. Gaussian Processes (GPs) offer a flexible and data-efficient foundation for statistical emulation, but their poor scalability limits applicability to large datasets. We introduce the Scaled Block Vecchia (SBV) algorithm for distributed GPU-based systems. SBV integrates the Scaled Vecchia approach for anisotropic input scaling with the Block Vecchia (BV) method to reduce computational and memory complexity while leveraging GPU acceleration techniques for efficient linear algebra operations. To the best of our knowledge, this is the first distributed implementation of any Vecchia-based GP variant. Our implementation employs MPI for inter-node parallelism and the MAGMA library for GPU-accelerated batched matrix computations. We demonstrate the scalability and efficiency of the proposed algorithm through experiments on synthetic and real-world workloads, including a 50M point simulation from a respiratory disease model. SBV achieves near-linear scalability on up to 512 A100 and GH200 GPUs, handles 2.56B points, and reduces energy use relative to exact GP solvers, establishing SBV as a scalable and energy-efficient framework for emulating large-scale scientific models on GPU-based distributed systems.
Batched Linear Algebra, Energy-Efficient HPC, Gaussian Processes, GPU Computing, High-Dimensional Data, Scalable Emulation, Vecchia Approximation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Many modern scientific and engineering advances rely on the use of computer models that use mathematical equations to represent physical systems and are implemented as high-fidelity simulations. These models allow researchers to explore phenomena that are impractical to observe experimentally, such as climate dynamics (Eyring et al., 2016; Kochkov et al., 2024), materials (Børvik et al., 2001; Walters et al., 2018), and cosmological structure formation (Springel, 2005; Spurio Mancini et al., 2022; Lawrence et al., 2017). Computer models can generate predictions that inform policy decisions, drive technological innovation, and facilitate fundamental discoveries by solving systems of differential equations or leveraging particle-based methods (Børvik et al., 2001; Eyring et al., 2016; Springel, 2005). However, when exploring models with many input parameters, learning the relationship between inputs and outputs may require thousands of evaluations (Fiore et al., 2018). Such high computational costs often constrain their usage, motivating the adoption of statistical emulators to approximate the simulator’s behavior at a fraction of the cost (Santner et al., 2003; Abdulah et al., 2024).
A common approach to building computer model emulators involves fitting a Gaussian Process (GP) to data generated by running the simulator at a predefined set of input values (design points), effectively learning a mapping from the model input space to the model output space (Andrianakis and Challenor, 2012). GP-based emulators enhance interpretability and facilitate efficient analysis, particularly when simulating scenarios across a broad parameter space. For example, variable importance (Linkletter et al., 2006; Saltelli et al., 2010) and sequential optimization (Winkel et al., 2021) may only be feasible using a cheap-to-evaluate GP emulator of the underlying expensive-to-evaluate simulator.
However, GPs face significant scalability challenges due to their $\mathcal{O}(n^{3})$ computational cost and $\mathcal{O}(n^{2})$ memory requirements, where $n$ is the number of data points. These costs arise from operations on a dense $n\times n$ covariance matrix that captures the correlations between data inputs  (Abdulah et al., 2018a). For instance, modeling 200200K locations demands about $300$ GB of memory and an estimated $2.6$ PFLOP of computation, making standard GPs infeasible for large-scale applications without approximation methods. Thus, numerous studies have addressed the computational and memory challenges of large-scale GPs. Some efforts have pursued exact GP computation on High-Performance Computing (HPC) systems (Abdulah et al., 2018a, 2019, 2021). For instance, the largest reported problem size was processed on ORNL’s Frontier system, handling 27.2427.24M data points in mixed-precision across approximately 99K nodes ($\sim$36K AMD GPUs) of the system (Abdulah et al., 2024). However, most of the literature has focused on approximation to alleviate the computational burden of GP modeling, including low-rank (Furrer et al., 2006; Bevilacqua et al., 2016; Shi et al., 2025), tile low-rank (Abdulah et al., 2018b; Salvaña et al., 2022; Mondal et al., 2023), covariance tapering (Furrer et al., 2006), and Vecchia approximations (Katzfuss and Guinness, 2021; Pan et al., 2024).
Among these approximation methods, Classic Vecchia (CV) approximation demonstrates superior advantages (Hazra et al., 2024). In (Katzfuss et al., 2022), Katzfuss et al. introduced the Vecchia approximation for GP emulation in the context of computer experiments and proposed the Scaled Vecchia (SV) method to enhance approximation accuracy in high-dimensional settings. However, their study was limited to small-scale problems and relied entirely on CPU-based computations. In contrast, efforts have been made to accelerate the classic Vecchia algorithm on GPUs, as seen in (Pan et al., 2024), which subsequently introduced the Block Vecchia (BV) approach (Pan et al., 2025) to enhance computational speed further while preserving approximation accuracy. Nevertheless, the use of CV approximation for high-dimensional GP emulation on GPUs remains unexplored. This gap presents an opportunity to develop a scalable GPU-based method that can efficiently address the computational and memory challenges posed by high-dimensional inputs.
In this work, we propose a novel algorithm called the Scaled Block Vecchia (SBV) approximation for distributed GPU-based systems. SBV integrates the Block Vecchia (BV) algorithm from (Pan et al., 2025) with the Scaled Vecchia approach introduced in (Katzfuss et al., 2022), while leveraging GPU acceleration techniques from (Pan et al., 2024). To the best of our knowledge, this work represents the first distributed implementation of any variant of the Vecchia-based GP algorithm. Leveraging the Message Passing Interface (MPI), we parallelize computations across multiple GPU-enabled nodes, enabling scalable performance well beyond the capabilities demonstrated in prior studies.
To validate the accuracy and robustness of our algorithm, we perform GP simulations across a range of parameter settings and benchmark its performance on the satellite drag dataset, a standard high-dimensional testbed for evaluating GP-based models (Sun et al., 2019). As an application in computer model emulation, we apply our algorithm to MetaRVM, a compartmental simulation model designed for generic respiratory virus diseases (Fadikar et al., 2025). Experimental results demonstrate both high predictive accuracy and strong scalability, with successful runs on up to 512 A100/GH200 GPUs and problem sizes reaching 2.562.56B data points. Additionally, our approach achieves a noticeable reduction in power consumption compared to the state-of-the-art exact GP framework on GPUs, i.e., ExaGeoStat (Abdulah et al., 2018a).
The paper is organized as follows: Section 2 summarizes the main contributions of the study. Section 3 provides a brief overview of related work. Section 4 presents the necessary background. Section 5 details the proposed algorithm. Section 6 describes the simulation study and reports the results of the computer model application. Section 7 focuses on performance evaluation and analysis. Finally, conclusions are presented in Section 8.
2. Contributions
The key contributions of this work are as follows:
  • We propose the Scaled Block Vecchia (SBV) approximation, a novel GP algorithm that integrates anisotropic input scaling with block-based conditioning to reduce both computational and memory complexity. SBV improves approximation accuracy in high-dimensional GPs.
  • We develop the first distributed implementation of any Vecchia-based GP variant designed for modern GPU-accelerated HPC systems. Our implementation employs MPI for inter-node parallelism and the MAGMA library (Abdelfattah et al., 2021; Agullo et al., 2009) for fine-grained batched GPU-accelerated linear algebra operations.
  • We benchmark SBV on synthetic datasets and the high-dimensional satellite drag dataset, demonstrating substantial improvements in runtime and predictive accuracy over state-of-the-art GP approximation methods.
  • We use and assess SBV to emulate the MetaRVM epidemiological simulation using 5050M samples and 10-dimensional inputs, achieving high predictive accuracy and practical scalability.
  • We demonstrate both weak and strong scaling on the JURECA-DC GPU (AMD EPYC + A100) and JUPITER (GH200 Superchip) systems at Jülich Supercomputing Centre (JSC), with experiments scaling up to $512$ nodes and problem sizes reaching up to 2.562.56B data points.
  • We perform a comprehensive energy analysis across GPU platforms, showing that SBV significantly reduces power consumption compared to exact GP frameworks such as ExaGeoStat while maintaining high throughput and accuracy.
3. Related Work
GP emulators: GPs are widely used for emulating high-fidelity simulators in scientific computing, with applications in cosmology (Spurio Mancini et al., 2022), climate modeling (Abdulah et al., 2024), and mechanism science (Sun et al., 2019). Notable examples include emulation of N-body cosmological simulations (Lawrence et al., 2017), particle-material interactions (Walters et al., 2018), and epidemiological models of disease spread (Fadikar et al., 2018). While recent advances in scalable GPs have improved performance, few approaches offer distributed, multi-GPU implementations. The proposed SBV algorithm directly addresses this gap by enabling efficient, scalable GP emulation on distributed, GPU-accelerated HPC systems.
CV approximation: CV approximation  (Vecchia, 1988; Katzfuss and Guinness, 2021) reduces the computational complexity of GPs by factorizing the joint distribution into a product of conditional distributions. Recent advancements have enhanced the effectiveness of the CV approximation through improved data point ordering and neighbor selection (Guinness, 2018). The BV approximation (Pan et al., 2025) builds on this by conditioning on blocks of observations rather than individual points, enabling batched linear algebra operations on GPUs and improving throughput and memory efficiency in high-performance environments (Pan et al., 2024). More recently, the SV method was introduced to address challenges associated with high-dimensional inputs by applying anisotropic input scaling (Katzfuss et al., 2022), which enables the model to capture directional-varying correlations. Our work extends these developments by integrating block-based conditioning and anisotropic scaling into a unified framework, i.e., SBV, and presenting a distributed implementation optimized for multi-GPU systems.
4. Background
This section provides background on GPs, the BV approximation, Kullback–Leibler (KL) divergence, and the scaled kernel function, which together form the foundation for our proposed extension for high-dimensional data. Table 1 summarizes the abbreviations used throughout the paper to enhance clarity and readability.
Table 1. List of abbreviations and their explanations.
Abbreviation Explanation
GPs Gaussian Processes
CV/SV/BV/SBV Classic/Scaled/Block/Scaled Block Vecchia
$nn^{*}$ Problem size in estimation/prediction
$bs_{est}$$bc_{est}$ ($bs/bc$) Block (batched) size/count for estimation
$bs_{pred}$$bc_{pred}$ ($bs^{*}/bc^{*}$) Block (batched) size/count for prediction
$m_{\text{est}}$$m_{\text{pred}}$ # nearest neighbors for estimation/prediction
4.1. Gaussian Processes (GPs)
GPs provide a flexible framework for modeling and predicting functions across low- and high-dimensional input spaces. Let $\bm{X}=[\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{n}]^{\top}$, where each $\bm{x}_{i}\in\mathbb{R}^{d}$, denote the $n$ input points with corresponding observations $\bm{y}=[y_{1},y_{2},\dots,y_{n}]^{\top}$. A GP with zero mean and kernel $K_{\bm{\theta}}(\bm{x}_{i},\bm{x}_{j})$ is formulated as $\bm{y}\sim\mathcal{N}(\bm{0},\Sigma_{\bm{\theta}})$ and its log-likelihood can be represented as shown in (Rasmussen and Williams, 2006):
(1) $\ell(\bm{\theta};\bm{y})=-\frac{n}{2}\log(2\pi)-\frac{1}{2}\log|\Sigma_{\bm{\theta}}|-\frac{1}{2}\bm{y}^{\top}\Sigma_{\bm{\theta}}^{-1}\bm{y},$
where $|\cdot|$ means determinant, $\Sigma_{\bm{\theta}}\in\mathbb{R}^{n\times n}$ has entries $\Sigma_{\bm{\theta}}]_{ij}=K_{\bm{\theta}}(\bm{x}_{i},\bm{x}_{j})$ parameterized by $\bm{\theta}$. For a set of $n^{*}$ new input points $\bm{X}^{*}=[\bm{x}_{1}^{*},\bm{x}_{2}^{*},\dots,\bm{x}_{n^{*}}^{*}]^{\top}$, we are interested in predicting their corresponding observations $\bm{y}^{*}=[y_{1}^{*},y_{2}^{*},\dots,y_{n^{*}}^{*}]^{\top}$. The joint distribution of $\bm{y}$ and $\bm{y}^{*}$ is assumed to be:
[…]
where $\Sigma_{\bm{\theta},*}\in\mathbb{R}^{n\times n^{*}}$ contains the covariance matrix between $\bm{X}$ and $\bm{X}^{*}$, and $\Sigma_{\bm{\theta}}^{*}\in\mathbb{R}^{n^{*}\times n^{*}}$ contains the covariances among $\bm{X}^{*}$. The conditional distribution of $\bm{y}^{*}$ given $\bm{y}$ is:
$p(\bm{y}^{*}|\bm{y})\sim\mathcal{N}(\mu^{*},\Sigma^{*}),$
where the conditional mean $\mu^{*}$ and covariance $\Sigma^{*}$ are derived as $\mu^{*}=\Sigma_{\bm{\theta},*}^{\top}\Sigma_{\bm{\theta}}^{-1}\bm{y},$ $\Sigma^{*}=\Sigma_{\bm{\theta}}^{*}-\Sigma_{\bm{\theta},*}^{\top}\Sigma_{\bm{\theta}}^{-1}\Sigma_{\bm{\theta},*}.$ The GPs have computational complexity $\mathcal{O}(n^{3})$ and memory complexity $\mathcal{O}(n^{2})$, making it infeasible on large scale problems.
4.2. Block Vecchia Method
The BV approximation (Pan et al., 2024) reduces the computational and memory cost of GPs by partitioning/clustering the dataset into $bc$ disjoint blocks, $\{\bm{B}_{1},\bm{B}_{2},\dots,\bm{B}_{bc}\}$. After applying a permutation $\zeta$ to order the blocks, the exact likelihood of $\bm{y}$ becomes:
$p_{\bm{\theta}}(\bm{y})=\prod_{i=1}^{bc}p_{\bm{\theta}}(\bm{y}_{\bm{B}_{i}^{\zeta}}|\bm{y}_{\bm{B}_{1}^{\zeta}},\bm{y}_{\bm{B}_{2}^{\zeta}},\dots,\bm{y}_{\bm{B}_{i-1}^{\zeta}}).$
The BV method approximates each conditional distribution by replacing each vector with a subset of neighbors $\bm{NN}(\bm{B}_{i}^{\zeta})$ by:
(2) $p_{\bm{\theta}}(\bm{y})\approx\prod_{i=1}^{bc}p_{\bm{\theta}}(\bm{y}_{\bm{B}_{i}^{\zeta}}|\bm{y}_{\bm{NN}(\bm{B}_{i}^{\zeta})}),$
where the number of $\bm{NN}(\bm{B}_{i}^{\zeta})$ is $m$ and $m\ll n$. The prediction for $\bm{y}^{*}$ under the BV framework approximates the conditional distribution $p(\bm{y}^{*}|\bm{y})$ by the product of block-level conditional distributions. Given the new blocks $\{\bm{B}_{1}^{*},\dots,\bm{B}_{bc^{*}}^{*}\}$ on $\bm{X}^{*}$, the predictive distribution is:
(3) $p_{\bm{\theta}}(\bm{y}^{*}|\bm{y})\approx\prod_{j=1}^{bc^{*}}p_{\bm{\theta}}(\bm{y}_{\bm{B}_{j}^{*}}|\bm{y}_{\bm{NN}(\bm{B}_{j}^{*})}),$
where $\bm{NN}(\bm{B}_{j}^{*})$ is the neighbors selected from the $\bm{y}$. The BV method reduces the computational complexity of the exact GPs from $\mathcal{O}(n^{3})$ to $\mathcal{O}(bc\cdot m^{3})$ and memory complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(bc\cdot m^{2})$. Figure 1 shows the computations involved in the BV algorithm (Pan et al., 2024), which includes three components, clustering $\{\bm{B}_{1},\bm{B}_{2},\dots,\bm{B}_{bc}\}$, Nearest Neighbor Search (NNS) $\bm{NN}(\bm{B}_{i}^{\zeta})$, and batched GPU computing.
Figure 1. The BV algorithm: (1) disjoint clustering, (2) Nearest Neighbor Searching (NNS), and (3) batched GPU log-likelihoods computation.
4.3. Kullback-Leibler (KL) Divergence
The KL divergence measures how one probability distribution diverges from a second probability distribution (Cover and Thomas, 2006). For two Gaussian distributions $\mathcal{N}_{0}$ and $\mathcal{N}_{1}$ with zero mean and covariance matrices $\Sigma_{0}$ and $\Sigma_{1}$, the KL divergence is (Murphy, 2012):
$D_{KL}(\mathcal{N}_{0}\|\mathcal{N}_{1})=\frac{1}{2}\left\{\text{tr}(\Sigma_{1}^{-1}\Sigma_{0})-n+\log\frac{|\Sigma_{1}|}{|\Sigma_{0}|}\right\}.$
For Vecchia-based GPs, previous work (Pan et al., 2025, 2024) shows that the KL divergence simplifies to the difference between exact and approximate log-likelihoods evaluated at $\bm{y}=\bm{0}$ as:
(4) $D_{KL}=\ell_{0}(\bm{\theta};\bm{0})-\ell_{a}(\bm{\theta};\bm{0}),$
where $\ell_{0}(\bm{\theta};\bm{0})$ is the exact log-likelihood and $\ell_{a}(\bm{\theta};\bm{0})$ is Vecchia-based approximated log-likelihood.
This work uses the KL divergence to quantify the discrepancy between the proposed SBV algorithm and posterior distributions from other GP methods.
4.4. Scaled Kernel Function
In high-dimensional spaces, covariance functions often require anisotropic scaling, where each input dimension is scaled according to its relevance to the output (Katzfuss et al., 2022). This allows the model to reflect varying sensitivities across dimensions and improves both accuracy and efficiency. The scaled covariance function is defined as
(5) $K_{\bm{\theta}}(\bm{x}_{k},\bm{x}_{k^{\prime}})=f\left(\left(\sum_{i=1}^{d}\frac{(x_{ki}-x_{k^{\prime}i})^{2}}{\beta_{i}^{2}}\right)^{\frac{1}{2}}\right),$
where $x_{ki}$ and $x_{k^{\prime}i}$ denote the $i$-th components of $\bm{x}_{k}$ and $\bm{x}_{k^{\prime}}$, and $\bm{\beta}=(\beta_{1},\dots,\beta_{d})^{\top}$ contains the dimension-specific range (scaling) parameters. $f(\cdot)$ is a kernel function, such as the Matérn kernel:
(6) $f(r)=\sigma^{2}\frac{2^{1-\nu}}{\Gamma(\nu)}r^{\nu}\mathcal{K}_{\nu}(r)+\sigma_{0}^{2}.$
Here, $r$ is the scaled distance and $\bm{\theta}=(\sigma^{2},\bm{\beta},\nu,\sigma_{0})^{\top}$ denotes the kernel parameters: $\sigma^{2}$ (process variance), $\sigma_{0}^{2}$ (nugget), and $\nu>0$ (smoothness). $\Gamma(\cdot)$ is the gamma function, and $\mathcal{K}_{\nu}(\cdot)$ is the modified Bessel function of the second kind of order $\nu$. Additional valid covariance classes are detailed in (Genton, 2001). We use the scaled form in Equation (5) to capture directional relevance, useful in emulation problems where only a subset of input dimensions significantly affect the output.
5. Distributed Scaled Block Vecchia Algorithm
This section presents the proposed distributed Scaled Block Vecchia (SBV) algorithm, consisting of data preprocessing and batched GPU-based log-likelihood computation. We then analyze the computational complexity of the algorithm, highlighting efficient GPU utilization.
5.1. GP Modeling
The distributed SBV algorithm consists of four key components: scaling and partitioning, Random Anchor Clustering (RAC), filtered subset selection for NNS, and batched GPU computation for optimizing the log-likelihood. The preprocessing steps, i.e., scaling and partitioning, RAC, and filtered NNS, are performed only once and executed on the CPU to prepare the desired data structure. The resulting structured data is then transferred to the GPU, allowing for the efficient execution of hundreds of parameter optimization iterations without redundant data movement. The distributed SBV algorithm is detailed in Algorithm 1, and the pipeline is illustrated in Figure 2, where Steps (2), (3), and (4) are the Block Vecchia (BV) algorithm pipeline.
Algorithm 1 Distributed SBV approximation algorithm
1:Input: Data $\{\bm{x}_{i},y_{i}\}_{i=1}^{n}$, dimension $d$, total block count $K$, local clusers $k_{p}$, nearest neighbors $m_{\text{est}}$, workers $P$, covariance function $K_{\theta}$, scaling parameters $\{\beta_{i}\}_{i=1}^{d}$.
2:Output: Approximate the log-likelihood $\ell$.
3:Step 0: Data Loading $X^{org}_{p}$
4: Each worker load local points $X^{org}_{p}$, $p=1,2,\ldots,P$.
5:Step 1: Partitioning and Scaling $X^{org}_{p}$
[…]
7:Step 2: Random Anchor Clustering $\bm{B}_{p}$
[…]
9: Randomly reorder these blocks $B_{p,i}$;
10:Step 3: NNS $\bm{J}_{p}$
[…]
12:Step 4: Batched Log-Likelihood Calculation $\log\mathcal{L}_{p}$
13:for each block $\bm{B}_{p},p=1,2,\ldots,P$ in Parallel do
14:  $\ell_{p}=\ell_{llh}(\bm{B}_{p},\bm{J}_{p},\bm{y}_{p,\bm{B}},\bm{y}_{p,\bm{J}})$, where $\bm{y}_{p,\bm{B}}$, $\bm{y}_{p,\bm{J}}$ are the observations associated with $\bm{B}_{p}$ and $\bm{J}_{p}$ respectively.
15:end for
16:Step 5: Reduction of Log-Likelihood Across Workers
17: Use MPI_Allreduce to sum $\ell_{p}$ from each worker and obtain the approximated log-likelihood $\ell$.
18:Return: The approximated log-likelihood $\ell$.
Figure 2. Distributed SBV pipeline. The pipeline for the distributed SBV algorithm: (0) parallel data loading; (1) anisotropic scaling and partitioning; (2) Random Anchor Clustering (RAC) to form disjoint spatial blocks; (3) neighborhood filtering for NNS within a radius $r=\lambda$; and (4) batched GPU computation of blockwise conditional log-likelihoods.
5.1.1. Data Partitioning and Scaling
Each worker $p$ begins by loading its assigned data subset $\bm{X}_{p}$. Firstly, the points are partitioned into several segments along their most relevant dimension, as shown in line 4 of Algorithm 2. Here, the most relevant dimension is denoted by $d_{max}$, and MPI_Alltoall is used to redistribute the dataset $X_{1:P}$, where $x_{p,i,d_{max}}$ refers to the $d_{max}$-th component of the $i$-th data point on worker $p$. Secondly, input dimensions are rescaled using anisotropic parameters $\bm{\beta}=[\beta_{1},\ldots,\beta_{d}]$. This scaling transforms each dimension $x_{ij}:=x^{org}_{ij}/\beta_{j}$, ensuring that the relevance of different dimensions is appropriately reweighted (see line 7 in Algorithm 2).
Figure 2 gives an example of partitioning within the unit square $0,1]^{2}$. During data loading, each color represents the portion assigned to a specific worker; for example, the green points are read by Worker 1. In the scaling and partition, the input space is partitioned into four segments along the most relevant dimension, $\beta_{2}=0.3$. After that, the input space is rescaled from the unit square to the rectangle $(0,1.5)\times(0,0.3)$, indicating that the first dimension (width) is more relevant to the response than the second (height). This rescaling and partitioning strategy ensures balanced and uniform data distribution across workers in the algorithm, enhancing computational efficiency while preserving accuracy.
Algorithm 2 Partitioning and Scaling $\mathcal{PS}(X^{org}_{1:P})$
1:Input: Complete dataset $X^{org}_{1:P}$.
2:Output: Scaled and redistributed $X_{1:P}$. Relevant dimension $d_{max}=\arg\max_{i}\{1/\beta_{i},i=1,2,\dots,d\}.$
3:for each worker $p$ in parallel do
4:  MPI_Alltoall, send $\bm{x}^{org}_{p,i}$ to $q$th node if $\text{int}(\bm{x}^{org}_{p,i,d_{max}}\cdot P)=q$.
5:end for
6:for each worker $p$ in parallel do
7:  Dimension scaling, $X_{p,j}:=X^{org}_{p,j}/\beta_{j}$, $j=1,2,\ldots,d$.
8:end for
9:Return: Redistributed and scaled $X_{1:P}$
5.1.2. Random Anchor Clustering (RAC)
Next, small blocks $\bm{B}_{p}$ are created, defined as $\bm{B}_{p}:=\{\bm{B}_{p,1},\ldots,\bm{B}_{p,k_{p}}\}$, for each partition by RAC independently. The RAC method is detailed in Algorithm 3. Here, we replace the K-means clustering in block Vecchia GPs (Pan et al., 2025) with RAC to reduce the computational cost induced by K-means for extensive datasets while maintaining comparable approximation accuracy. The RAC algorithm begins by randomly selecting $k_{p}$ anchors (centers) for each worker $p=1,2,\dots,P$ in parallel. These anchors serve as the centers of blocks. Subsequently, each $x_{p,i}$ is assigned to the $j$th block based on their distance, as shown in line 5 of Algorithm 3, thus constructing clusters, $B_{p}:=(B_{p,1},B_{p,2},\ldots,B_{p,k_{p}})$ with $p=1,2,\dots,P$. The RAC method does not involve any communication between workers, considering that the size of clusters is typically much smaller than the overall dataset. For example, a single cluster may contain only 100, while the entire dataset consists of 10 million locations. In Step 2 of Figure 2, each worker creates 4 blocks in their data partition.
Algorithm 3 RAC algorithm $\mathcal{C}(X_{1:P})$
1:Input: $X_{1:P}$ full dataset.
2:Output: $B_{1:P}$ set of cluster for each node.
3:for each worker $p$ in parallel do
4:  Randomly choose $k_{p}$ local centers $C_{p}=\{\bm{c}_{p,i}\mid i=1,2,\dots,k_{p}\}$ in $X_{p}$, and initialize $B_{p,j}=\{\bm{c}_{p,i}\}.$
5:  Assign $\bm{x}_{p,i}$ to $j$th cluster $B_{p,j}$, where $j=\arg\min_{j}\|\bm{x}_{p,i}-\bm{c}_{p,j}\|^{2},j=1,2,\dots,k_{q}$.
6:end for
7:Return: Clusters $B_{1:P}$
5.1.3. Filtered $m$-Nearest Neighbor Search (NNS)
With block sets $\bm{B}_{1:P}$, the next goal is to find the exact $m$ nearest neighbors for all block centers in high-dimensional space, where search areas are varied for each query center in the Vecchia-based GPs, see Equation (2). Existing methods (Abbasifard et al., 2014; Ferhatosmanoglu et al., 2001; Clarkson et al., 2006) for exact $m$-NNS with query-specific search areas face limitations, especially for large datasets.
To address these issues, we adopt the filtered $m$-NNS on each block $B_{p,i}$, detailed in Algorithm 4. This approach selects a small filtered subset (a circle of radius $\lambda$) for brute-force $m$-NNS to obtain exact results while reducing computational cost. In Algorithms 1 and 4, $\mathcal{V}$ denotes the filtered $m$-NNS function, which begins by computing a Monte Carlo-based distance threshold. This threshold $\lambda$ is calculated using:
(7) $\lambda=\left(\alpha\frac{m\zeta}{n}\right)^{1/d},\quad\zeta=\begin{cases}\frac{\Gamma\left(\frac{d}{2}+1\right)}{\pi^{\frac{d}{2}}}&d\text{ is even}\\ \frac{2\pi^{\frac{d-1}{2}}\Gamma\left(\frac{d+1}{2}\right)}{\Gamma(d+1)}&d\text{ is odd}\end{cases}.$
Here, $n$ is the total number of points, $m$ is the number of nearest neighbors, $d$ is the dimensionality, and $\alpha$ is an expansion factor to account for irregularities in point distribution (e.g., $\alpha=100$ implies the candidate set is expected to be 100 times larger than $m$). Then, we prepare the small filtered subset, which consists of two steps, coarser and finer preparation. In the coarser preparation, each worker redistributes candidate clusters to the corresponding worker if the distance between two centers is within the calculated threshold using MPI_Alltoall. The complementary coarser candidate set is denoted as $B_{p}^{cand}$ (lines 10-13 in Algorithm 4). In the finer partition, each worker independently selects finer candidates $S_{p,i}$ for the $i$th cluster, i.e., $\bm{x}_{k}$ is added to the finer candidate set $S_{p,i}$ if the distance between the $i$th local center and the data point $\bm{x}_{k}$ is within the threshold $\lambda$ (lines 15-22 in Algorithm 4). Finally, each local center $\bm{c}_{i}$ applies brute-force search for its $m$ nearest neighbors $J_{p,i}$ from the finer candidates set $S_{p,i}$. The computational cost of brute search in $S_{p,i}$ is negligible compared to $\mathcal{O}(n)$ due to $\alpha m\ll n$ and total block count $K\ll n$, thus filtered $m$-NNS achieves computational complexity $\mathcal{O}(\alpha m+K)$ and has a simple implementation.
Figure 3 illustrates the filtered m-NNS using an example of four workers. In the coarser candidate preparation, MPI_Alltoall is utilized to distribute related blocks (within distance threshold $\lambda$) to the 33rd worker, forming the coarse candidates $B_{3}^{cand}$. Subsequently, the 33rd block $B_{3,3}$ refines candidates for the brute $m$-NNS (within distance threshold $\lambda$). Finally, blocks with their nearest neighbors and their observations are constructed as a contiguous arrangement for batched GPU computing.
Algorithm 4 NNS algorithm based on filtered subset $\mathcal{V}(\bm{B}_{p},m)$
1:Input: $\bm{B}_{p}$ cluster set, nearest neighbors $m$, expansion factor $\alpha=100$ (default).
2:Output: Nearest neighbor set $J_{1:P}$
3:Step1: Distance Threshold and Centers
4: Calculate distance threshold, $\lambda$ according to Equation (7).
5:for each block $\bm{B}_{p,i}$ in $\bm{B}_{p}$ do
6:  Update centers $C_{i}=\{\bm{c}_{p,i}\}$, where $\bm{c}_{p,i}=\frac{1}{k_{p,i}}\sum_{i=1}^{k_{p,i}}\bm{x}_{p,i}$ with $\bm{x}_{p,i}\in\bm{B}_{p,i}.$
7:end for
8:MPI_Allgather, gather all centers $C_{1:P}$ to each worker.
9:Step 2: Prepare Coarser Candidates
10:for each worker $p$ in parallel do
11:  MPI_Alltoall, redistribute $\bm{B}_{p,i}$ to $q$th node if $\|\bm{c}_{p,i}-\bm{c}_{q,j}\|^{2}\leq\lambda$ where $j$ exists.
12:  Received clusters form the candidate set $\bm{B}^{cand}_{p}$ for worker $p$.
13:end for
14:Step 3: Prepare Finer Candidates for NNNN
15:for local block center $\bm{c}_{i}$ in $\{\bm{c}_{i}\}_{i=1}^{k_{p}}$ do
16:  for candidate block center $\bm{c}_{j}$ in $\bm{B}^{cand}_{p}$ do
17:   if The order $\bm{c}_{j}$ not exceeds order $\bm{c}_{i}$ then
18:    for each point $\bm{x}_{k}$ in $\bm{B}^{cand}_{p,j}$ do
19:     Calculate the distance $d(\bm{c}_{i},\bm{x}_{k})$. If $d(\bm{c}_{i},\bm{x}_{k})<\lambda$, add $\bm{x}_{k}$ to the candidate set $S_{p,i}$ for block $B_{p,i}$.
20:    end for
21:   end if
22:  end for
23:  Searching $m$ nearest neighbors for $i$th block, i.e., $\bm{J}_{p,i}=NNS(\bm{c}_{i},S_{p,i},m)$.
24:end for
25:Return: Nearest neighbor set $\{\bm{J}_{p},p=1,2,\ldots,P\}.$
Figure 3. Filtered NNS algorithm pipeline. (1) workers expand local partitions by radius $\lambda$ and exchange boundary data; (2) finer candidates are selected within $r=\lambda$ around each block; (3) $m$ nearest neighbors are identified; (4) data and covariance matrices are generated.
5.1.4. Batched Log-likelihood Computations
Once the data preprocessing is finished, we copy clusters $\bm{B}_{1:P}$, nearest neighbors $\bm{J}_{1:P}$, and their observations to GPU global memory for log-likelihood optimization. Algorithm 5 outlines the computation of a single iteration of log-likelihood, which needs to be evaluated multiple times until the prespecified optimization configuration is achieved. Here, the batched operations execute for log-likelihood computation. This involves constructing local covariance matrices $\Sigma^{lk},\Sigma^{cross}$ and $\Sigma^{con}$ using the scaled kernel $K_{\bm{\theta}}$. Subsequently, the GPU data structure is constructed in Step 4 in Figure 2. Following, batched Cholesky decompositions (POTRF) is utilized to factorize neighbor covariance matrices, $\Sigma^{con}$. Conditional mean and covariance updates are computed via batched triangular solver (TRSM) and matrix multiplications (GEMM), while batched triangular solver (TRSV) and determinant calculations derive per-block log-likelihoods $\ell_{p,i}$. Each node aggregates its local computations $\ell_{p}=\sum_{i=1}^{k_{p}}\ell_{p,i}$, and a global reduction across all nodes yields the final log-likelihood $\ell=\sum_{p=1}^{P}\ell_{p}$ using MPI_Allreduce.
Algorithm 5 Batched Block Vecchia $\ell_{llh}(B_{p},J_{p},\bm{y}_{p,B},\bm{y}_{p,J})$
1:Input: In $p$th node, $B_{p}$ cluster set, $J_{p}$ nearest neighbor set, $\bm{y}_{p,B}$ and $\bm{y}_{p,J}$ observation sets.
2:Output: log-likelihood $\ell_{p}$
3:Step 1: Conditional Mean and Covariance Update
[…]
14:Step 2: GPs calculations
[…]
20:return  log-likelihood $\ell_{p}$
5.1.5. GP Prediction
During the prediction stage, the pipeline follows the estimation process, with the key difference being that in Step (2) GP calculations in Algorithm 5 are replaced by conditional simulations. Specifically, the variance vector, $\bm{\sigma}=\left(\sigma_{1},\sigma_{2},\dots,\sigma_{n^{*}}\right)^{\top}$ is subtracted from the diagonal elements of the covariance matrices $\mathbf{\Sigma}^{new}$, and $\bm{\mu}^{new}$ serves as the predicted values, $({y}_{*1},{y}_{*1},\dots,{y}_{*n^{*}})$. Next, conditional simulations are performed by drawing 1000 samples from the normal distribution $\mathcal{N}(y_{*j},\sigma_{j})$. Then, the sample mean $\tilde{\mu}_{j}$ and variance $\tilde{\sigma}_{j}^{2}$ are computed, and the 95% confidence interval is given by, $(\tilde{\mu}_{j}-z_{\alpha/2}\tilde{\sigma}_{j},\tilde{\mu}_{j}+z_{\alpha/2}\tilde{\sigma}_{j}),$ where $\alpha=0.05,j=1,2,\dots,n^{*}$, which follows the same as (Pan et al., 2025).
Overall, the workflow minimizes inter-node communication and computations by scaling and partitioning, RAC, and filtered $m$-NNS, while GPU-batched kernels exploit fine-grained parallelism for covariance and log-likelihood computations. Moreover, this algorithm incorporates anisotropic scaling for high-dimensional inputs to ensure accurate emulation for computer models. This framework enables scalable and efficient log-likelihood evaluation for large-scale GPs in the high-dimensional problem setup.
5.2. Complexity Analysis
We analyze the memory and computational complexity of SV and SBV, focusing on the estimation of GPU stage, which dominates due to repeated iterations, unlike the one-time CPU preprocessing.
For SV, the memory footprint is $\mathcal{O}(nm^{2}+nm)$, primarily due to the storage of $n$ small covariance matrices $\bm{\Sigma}_{1:n}$ and corresponding conditioning vectors $\bm{y}_{1:n}$. In contrast, SBV partitions the data into $bc$ blocks. Each block stores three covariance matrices, i.e., $\bm{\Sigma}_{i}^{\text{lk}}$, $\bm{\Sigma}_{i}^{\text{cross}}$, and $\bm{\Sigma}_{i}^{\text{con}}$, and two observation vectors: $\bm{y}_{J_{i}}$ for the conditioning set and $\bm{y}_{B_{i}}$ for the block itself. Assuming an average block size of $bs=n/bc$, the memory complexity for each SBV component is: $\mathcal{O}(n\cdot bs)$ for $\bm{\Sigma}_{i}^{\text{lk}}$, $\mathcal{O}(mn)$ for $\bm{\Sigma}_{i}^{\text{cross}}$, $\mathcal{O}(bc\cdot m^{2})$ for $\bm{\Sigma}_{i}^{\text{con}}$, $\mathcal{O}(bc\cdot m)$ for $\bm{y}_{J_{i}}$, and $\mathcal{O}(n)$ for $\bm{y}_{B_{i}}$. Therefore, the total memory complexity is $\mathcal{O}(nm^{2}+nm)$ for SV and $\mathcal{O}(n\cdot bs+mn+bc\cdot m^{2})$ for SBV.
The dominant computational costs in SBV arise from Cholesky factorizations and matrix-matrix multiplications. The Cholesky operations contribute $\mathcal{O}(bc\cdot bs^{3})$ for block matrices and $\mathcal{O}(bc\cdot m^{3})$ for the conditioning sets. Matrix-matrix multiplications add a complexity of $\mathcal{O}(mn\cdot bs)$. Thus, the total computational complexity for SBV is $\mathcal{O}(n\cdot bs^{2}+bc\cdot m^{3}+mn\cdot bs)$, compared to $\mathcal{O}(nm^{3})$ for SV, which performs a Cholesky factorization for each data point.
For high-dimensional input spaces, we recommend setting the number of nearest neighbors to four times the block size, i.e., $m=4\times bs$. Under this setting, the complexities are simplified and summarized in Table 2. SBV achieves lower memory and computational costs than SV, especially when leveraging GPU-batched operations for large-scale GP models.
Table 2. Complexity analysis for SV and SBV under $m=4\times bs$.
Method Memory Computational
SV $\mathcal{O}(nm^{2})$ $\mathcal{O}(nm^{3})$
SBV $\mathcal{O}(nm)$ $\mathcal{O}(nm^{2})$
6. Simulation and Benchmark Evaluation
This section begins with a synthetic GP simulation using the Matérn kernel in high-dimensional input space, highlighting the critical role of scaling and clustering. We then apply the proposed SBV algorithm to the satellite drag benchmark dataset to demonstrate its effectiveness and accuracy in emulating complex computer models.
6.1. Synthetic Data Simulation
(a) KL divergence for fitting.
(b) Prediction accuracy (MSPE).
(c) block size (bsbs) and SBV accuracy.
Figure 4. Comparison of Vecchia-based GP methods on model fitting and prediction accuracy, i.e., Classic Vecchia (CV), Block Vecchia (BV), Scaled Vecchia (SV), Scaled Block Vecchia (SBV). Subfigures (a) and (b) evaluate the GP variants fitting (KL divergence) and prediction separately, showing that scaling and block enable better accuracy; subfigure (c) Investigate SBV with different block sizes (bs), showing that a larger $m$ and a smaller block size yield the best fitting accuracy.
The simulation framework follows the design outlined in (Katzfuss et al., 2022), ensuring consistency for comparative analysis. The synthetic dataset is generated from a zero-mean GP with a Matérn kernel ($\nu=3.5$) using Equation (6). The input space is, $\bm{x}\in[0,1]^{10}$. The response is modeled as $y(\mathbf{x})\sim\mathcal{N}(\bm{0},K_{\bm{\theta}}(\mathbf{x},\mathbf{x}^{\prime}))$, where the parameter vector is given by $\bm{\theta}=(\sigma^{2},\tau^{2},\bm{\beta})$. The true parameters are set as $(\sigma^{2},\tau^{2})=(1.0,0)$, with range parameters $\beta_{1}=\beta_{2}=0.05$ for the relevant dimensions and $\beta_{3}=\beta_{4}=\dots=\beta_{10}=5$ for the irrelevant ones. We evaluate four Vecchia-based GPs on this synthetic dataset: CV (Vecchia, 1988), BV (Pan et al., 2025), SV (Katzfuss et al., 2022), and our proposed SBV. The choice of $bs_{est}=bs_{pred}=10$ in BV and SBV reflects a balance between computational efficiency and predictive accuracy. Model performance is assessed using both the KL divergence defined in Equation (4) and the Mean Squared Prediction Error (MSPE). To isolate approximation error and avoid conflating it with parameter estimation error, the true parameters $(\sigma^{2},\tau^{2},\bm{\beta})$ are directly supplied to all four models.
The results are presented in Figure 4. In Figure 4a, the proposed SBV method achieves the lowest KL divergence, followed by SV, BV, and CV, highlighting the advantage of combining input scaling and clustering over the CV approach. The performance gains of SBV arise from its ability to address input anisotropy through joint scaling and to increase neighborhood coverage through clustering. Figure 4b shows a similar trend regarding predictive accuracy, where SBV again yields the lowest MSPE, followed by SV, BV, and CV. These results are consistent with the findings on the KL divergence and further validate the benefits of incorporating input scaling and structured clustering. In Figure 4c, we examine the impact of block size on KL divergence. Smaller blocks ($bs_{est}=12$) result in lower KL divergence compared to larger ones ($bs_{est}=50$), indicating improved approximation accuracy with finer partitioning. However, larger blocks impose heavier computational loads, which in turn allow for effective GPU utilization. Therefore, the choice of block size $bs$ introduces a trade-off between accuracy and computational cost.
6.2. Evaluation on Satellite Drag Benchmark
The satellite drag dataset in (Sun et al., 2019) is a widely used benchmark for evaluating GP-based models in high-dimensional settings (Sun et al., 2019; Katzfuss et al., 2022). It is generated from a high-fidelity simulator that models atmospheric drag coefficients in low Earth orbit (LEO). The dataset comprises 22M simulation runs for each of six primary atmospheric species: atomic oxygen ($O$), molecular oxygen ($O_{2}$), atomic nitrogen ($N$), molecular nitrogen ($N_{2}$), helium ($He$), and atomic hydrogen ($H$). The total drag coefficient is computed as a weighted average of these species. Each species-specific dataset has an 8-dimension input space: relative velocity, surface temperature, atmospheric temperature, yaw angle, pitch angle, and two accommodation coefficients.
Following the experimental design in (Katzfuss et al., 2022), we split the dataset into 90% for training and 10% for testing and performed 10-fold cross-validation for robust evaluation. We adopt the same Matérn kernel with nugget effects, and smoothness parameter $\nu=3.5$ as in (Katzfuss et al., 2022). Model performance is assessed using the Root Mean Squared Percentage Error (RMSPE) and the estimated kernel parameters. The models compared in this study are mentioned in Table 3.
Table 3. Vecchia-based GP configurations on the satellite drag dataset, showing block sizes ($bs_{\text{est}}$, $bs_{\text{pred}}$) and neighbor counts ($m_{\text{est}}$, $m_{\text{pred}}$) for estimation and prediction.
Model SV SBV1 SBV2 SBV3 SBV4 SBV5 SBV6
$bs_{\text{est}}$ 1 100 100 100 100 100 100
$bs_{\text{pred}}$ 1 5 5 5 5 5 5
$m_{\text{est}}$ 50 200 200 200 400 400 400
$m_{\text{pred}}$ 140 200 400 600 200 400 600
For model fitting, SV uses $m_{\text{est}}=50$ nearest neighbors on a 50K-sample subset, following the configuration in (Katzfuss et al., 2022), which reported only marginal gains in predictive accuracy beyond this setting relative to the increased computational cost. In contrast, SBV leverages the full dataset with larger neighbor counts ($m_{\text{est}}=200$ or $400$) while maintaining a consistent average block size of $bs_{\text{est}}=100$ across all configurations. Due to the high memory and computational demands of SV, fitting the full dataset on a single GPU is infeasible, making SBV a more practical and scalable alternative.
Figure 5 presents the RMSPE results for Vecchia-based GP models on the benchmark dataset. SBV consistently achieves the lowest RMSPE, demonstrating the effectiveness of combining input scaling with clustering. Increasing the number of prediction neighbors $m_{\text{pred}}$ improves accuracy; for example, SBV with $m_{\text{pred}}=400$ reduces RMSPE by approximately 15% on average compared to $m_{\text{pred}}=200$. To better understand the improved accuracy achieved by SBV, Figure 6 compares the estimated inverse lengthscales ($1/\beta_{i}$) for SV and SBV. Both methods identify the final three input dimensions as most relevant while treating the others as less influential. However, the parameter estimates produced by SBV differ significantly from those of SV, underscoring the importance of using a larger number of nearest neighbors for accurate GP fitting. From a computational and memory perspective, SV faces challenges in handling large neighbor counts due to substantial cost and memory usage increases, which limit efficient GPU utilization. In contrast, SBV remains scalable and efficient under these conditions.
Figure 5. Root Mean Squared Percentage Error (RMSPE) of Vecchia-based GP models on six outputs from the satellite drag benchmark.
Figure 6. Estimated input relevance ($1/\beta_{i}$) for SV and SBV on the satellite drag benchmark.
6.3. MetaRVM Emulator
MetaRVM (Fadikar et al., 2025) is an R package that simulates the spread of respiratory viruses using a graph-based probabilistic model. By dividing a total population into sub-populations such as zones, age groups, or races, it models disease dynamics within and between these groups. This structure enables the simulation of interactions and disease spread within and between subpopulations. In healthcare, such models are crucial for understanding transmission patterns, evaluating intervention strategies, and informing public health policies. They help predict outbreak trajectories, assess the potential impact of vaccinations, and optimize resource allocation, ultimately aiding in the control and prevention of respiratory infections. The parameters in MetaRVM simulator are listed in Table 4. To understand the relationship between the input parameters and the accumulated number of hospitalizations (output) over 100 days in one population, we randomly choose 50M sets of input variables and obtain the accumulated hospitalizations using MetaRVM simulator. The 50M outputs are then divided into 90% for fitting SBV GPs with Matérn kernel at $\nu=3.5$ and 10% is used for prediction, where the RMSPE is reported. The input is also scaled into $[0,1]$, and the output is normalized with a mean of 1 to avoid the abnormal values in RMSPE. In the SBV, we set $bs_{\text{est}}=100,bs_{\text{pred}}=25$ and investigate the accuracy as increasing the number of nearest neighbors.
Table 4. The MetaRVM dataset simulation parameters.
Input Meaning Bound
tsts transmissibility for susceptible (0.1, 0.9)
tvtv transmissibility for vaccinated (0.1, 0.9)
dvdv mean duration in vaccinated state (30, 90)
dede mean duration in exposed state (1, 5)
dpdp mean duration in infectious presymptomatic state (1, 3)
dada mean duration in infectious asymptomatic state (1, 9)
dsds mean duration in infectious symptomatic state (1, 9)
dhdh mean duration in hospitalized state (1, 5)
drdr mean duration in recovered state (30, 90)
veve vaccine efficacy (0.3, 0.8)
The results are presented in Figure 7, where a large $m$ for both estimation and prediction generally leads to improved RMSPE. In the parameter estimation, the relevance $dh$ and $dr$ is close to 0, which aligns with our expectation, as these parameters are not involved in the accumulated number of hospitalizations in the MetaRVM simulator. Besides, for $ds$ and $tv$, the estimation varies significantly to the increasing value of $m_{\text{est}}$.
(a) RMSPE.
(b) Estimated parameters.
Figure 7. Root Mean Squared Percentage Error (RMSPE) and estimated relevance of SBV algorithm on MetaRVM dataset.
7. Performance and Energy Evaluation
(a) Runtime on AMD EPYC/ NVIDIA A100 GPU (CPU+GPU).
(b) Performance on AMD EPYC/ NVIDIA A100 GPU (GPU only).
(c) Runtime on GH200 superchip (CPU+GPU).
(d) Performance on GH200 superchip (GPU only).
Figure 8. Performance comparison of SBV and SV methods with 500 MLE iterations on single AMD EPYC with NVIDIA A100 (40 GB) and single NVIDIA GH200 superchip (96 GB).
We conduct experiments on two modern NVIDIA GPU architectures, A100 and GH200, hosted at the Jülich Supercomputing Centre (JSC) on the JURECA-DC and JUPITER systems including the early access JUPITER development system JEDI. JUPITER features about 6000 nodes, each equipped with four NVIDIA GH200 Grace Hopper Superchips. 48 of these nodes were accessible early as the JEDI system. Each Superchip integrates a 72-core Grace CPU (3.1 GHz, 120 GB) and a Hopper GPU (96 GB HBM3), connected via NVLink-C2C (900 GB/s), with a total power consumption of approximately 680 W per Superchip. JURECA-DC contains 192 GPU-equipped nodes, each with two 64-core AMD EPYC 7742 CPUs (2.25 GHz, 512 GB, 225 W per socket) and four NVIDIA A100 GPUs (40 GB HBM2, 400 W per GPU). Both systems use InfiniBand interconnects: JUPITER and JEDI provide four NDR200 (200 Gbit/s) links per node, while JURECA-DC GPU offers two HDR (200 Gbit/s) links per node. Our framework is compiled with GCC 13.3.0, uses CUDA 11.8/12.0, and links against MAGMA 2.7.2 (Abdelfattah et al., 2021; Agullo et al., 2009) and NLopt 2.7.1. All experiments run in double precision and are repeated for consistency.
All following experiments set the default $bs=100$, 10-dimensional input $0,1]^{10}$, and only consider the estimation stage, which is the most time-consuming part in the whole lifetime of emulations; thus, $m$ represents the number of nearest neighbors for estimations and uses Maximum Likelihood Estimation (MLE). Here, we set the iterations to 500, considering that the average number of optimization rounds is 500 in benchmark datasets.
7.1. Single-Node Performance
Figure 8 illustrates the runtime and GPU throughput performance of the SBV and SV methods on two nodes: an NVIDIA A100 with its host CPU, and the GH200 Superchip. Subfigures (a) and (c) show that SBV consistently outperforms SV in total execution time across different values of $m$, demonstrating superior scalability as the dataset size increases. As expected, larger values of $m$ increase the total runtime due to higher computational cost per iteration; however, they also lead to significantly better approximation and prediction accuracy, as shown in the accuracy section.
Subfigures (b) and (d) highlight the GPU throughput (in GFLOP/s) for a single iteration, showing that SBV achieves significantly higher sustained performance than SV on both platforms. Notably, the GH200 delivers performance exceeding $1600$ GFLOP/s with $m=400$. These results confirm that SBV is faster and more GPU-efficient than SV, particularly when utilizing larger neighborhoods on modern architectures, such as the GH200.
7.2. Scalability Across Multiple Nodes
Figure 9 demonstrates the weak and strong scaling performance of the distributed SBV implementation on NVIDIA A100 and GH200 architectures, scaling up to 512 GPUs. In weak scaling (Subfigures (a) and (c)), both systems maintain high parallel efficiency ($PE$) as the problem size increases proportionally with the number of GPUs. The drops in the larger scale are derived from the CPU computation, specifically Step 2 in the Algorithm 4, which can be reduced using a smaller $\lambda$ in Eq 7. Specifically, the candidate count does not scale proportionally and retains some stochastic variability, introducing non-uniform load balancing across tasks. A gradually decreasing $\lambda$ is recommended in the NNS as the GPU count increases.
In strong scaling (Subfigures (b) and (d)), total execution time decreases nearly linearly with increasing GPU count, and parallel efficiency exceeds ideal scaling ($PE>1.0$) in several cases due to less nearest neighbor candidates in the Algorithm 4, where $\lambda$ can be increased accordingly to ensure there is enough candidate for NNS. The drops in the larger GPU count originate from the overhead of multiple function calls, where the 500 iterations are executed and their accumulation cannot be hidden as the computational task is light for each GPU, where each iteration only costs around 30 milliseconds.
In both scaling experiments, the influence of neighborhood size $m_{\text{est}}$ is also evident: larger values, such as $m=400$, increase the computational workload per point and runtime, but result in better GPU throughput and overall hardware utilization. Smaller values, such as $m=100$, result in faster runtimes but less efficient GPU utilization, especially on the GH200. Mid-range values (e.g., $m=200$) balance performance and efficiency.
A detailed breakdown shows that GPU compute time dominates execution, accounting for over 80% of the total runtime in large problem sizes. However, the remaining CPU-side computation introduces variations in scaling. On A100, limited CPU resources (32 cores per task) contribute to a gradual drop in $PE$, whereas the 72-core CPU on GH200 helps better hide this overhead. This difference comes from the role of the CPU in maintaining a globally sorted array for block reordering and candidate filtering, as in Equation (2).
(a) Weak scalability and PE (A100) (2M–1024M points).
(b) Strong Scalability and PE (A100) (2M points).
(c) Weak scalability and PE (GH200) (5M–2.56B points).
(d) Strong Scalability and PE (GH200) (5M points).
Figure 9. Weak and strong scaling of SBV were evaluated on up to 512 GPUs, using 500 MLE iterations on two architectures: AMD EPYC with NVIDIA A100 (40 GB, up to 128M points) and the NVIDIA GH200 superchip (96 GB, up to 2.56B points).
7.3. Power Consumption Analysis
(a) Single NVIDIA A100 GPU (2M points).
(b) Single NVIDIA H100 GPU (5M points).
Figure 10. Power consumption/energy (kJ) on a single GPU for two NVIDIA GPUs over 500 iterations per $m_{\text{est}}$. A100 GPU with 400 W max power cap and Hopper GPU with (680 W - the power usage of the Grace CPU, RAM, and system I/O) max power cap.
Figure 10 presents the power consumption profiles of a single A100 and GH200 superchip GPU while running 500 log-likelihood iterations across varying neighborhood sizes ($m_{\text{est}}$), using a problem size of 2M points on A100 and 5M on GH200, the largest size that fits in GPU memory. As expected, both power draw and total energy consumption increase with larger $m_{\text{est}}$ due to the higher computational workload. On A100 (Figure 10a), power remains well below the 500 W limit, with total energy ranging from 42.7 kJ ($m_{\text{est}}=100$) to 148.6 kJ ($m_{\text{est}}=400$). In contrast, the Hopper GPUs as part of the GH200 superchips (Figure 10b) consume significantly more energy, up to 343.4 kj, while handling a larger problem size, enabled by a higher power envelope. For GH200 superchips, the maximum power consumption is per superchip (680 W for each GH200 in JEDI) and the Grace CPU is prioritized. To avoid power shortage and thus performance drops of the Hopper GPUs, the maximum CPU power consumption can be capped (including RAM and system I/O). On the JEDI system, the default CPU power cap is 100 W, leaving at least 580 W for the Hopper GPUs.
[…]
8. Conclusion
This paper introduced the Scaled Block Vecchia (SBV) approximation, a novel framework for scalable GP emulation tailored for high-dimensional datasets on distributed GPU systems. By integrating anisotropic input scaling with block-based conditioning and leveraging batched GPU kernels, SBV substantially reduces memory and computational costs while maintaining high predictive accuracy. To our knowledge, this is the first distributed implementation of a Vecchia-based GP approximation. Through extensive experiments on synthetic and real-world data, including the satellite drag benchmark and MetaRVM epidemiological model, we demonstrate the scalability up to 512 GPUs and the ability to handle up to 2.52.5B points with energy efficiency and state-of-the-art accuracy.
Acknowledgments
This work was supported and funded by King Abdullah University of Science and Technology (KAUST). We gratefully acknowledge Ahmad Abdelfattah from the University of Tennessee, Knoxville (UTK), for his support and assistance in optimizing the use of the MAGMA library. We also thank the KAUST Supercomputing Laboratory (KSL) team for providing access to the Ibex cluster for initial runs and experiments on available GPUs. We further acknowledge the computing time provided by the John von Neumann Institute for Computing (NIC) on the supercomputer JURECA-DC. We are grateful for access to JEDI and JUPITER, which is provided through the JUPITER Research and Early Access Program (JUREAP).